Write the half-reactions as they occur at each electrode and the net cell reaction for this electrochemical cell containing copper and silver: Cu(s)|Cu2 (aq)||Ag (aq)|Ag(s)

Anode?

Cathode?

Net cell reaction?
Answer #1
Concepts and reason

Deduce the metals present at the cathode and anode. Write the oxidation half-cell reaction for the metal present at anode and reduction half-cell reaction for the metal present at cathode. Balance the electrons on both the reactions.

Add the reactions to get the net cell reaction.

Fundamentals

Oxidation is defined as the process in which a substance loses electrons and gets oxidized whereas reduction is defined as the process in which the substance accepts electrons and gets reduced.

Electrochemical cell is a chemical cell which generates electrical energy from the chemical reaction taking place in the reaction. It consists of cathode and anode.

Cathode is defined as electrode through which the electrons enter the cell. It is negatively charged electrode. In this electrode, the reduction process takes place.

Anode is defined as electrode through which the electrons leave the cell. It is positively charged electrode. In this electrode, the oxidation process takes place.

Cell notation is a shorthand notation which is used for denoting the reactions taking place in the electrochemical cell. The cells denoting anode and cathode are separated by vertical bars, which represents a salt bridge.

The general cell notation is written as follows:

X(s)X+(aq)Y+(aq)Y(s)AnodeCathode\begin{array}{l}\\{\rm{X}}\left( s \right)|\,{{\rm{X}}^ + }\left( {aq} \right)\,||\,{{\rm{Y}}^ + }\left( {aq} \right)\,|\,{\rm{Y}}\left( s \right)\\\\\,\,\,\,\,\,{\rm{Anode}}\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{Cathode}}\\\end{array}

The anode and cathode for the given cell notation are as follows:

Anode: Copper

Cathode: Silver

The reaction taking place at anode, that is, oxidation, is written as follows:

Cu(s)Cu2+(aq)+2e{\rm{Cu}}\left( s \right) \to {\rm{C}}{{\rm{u}}^{2 + }}\left( {aq} \right) + 2{{\rm{e}}^ - }

The reaction taking place at cathode, that is, reduction, is written as follows:

Ag+(aq)+eAg(s){\rm{A}}{{\rm{g}}^ + }\left( {aq} \right) + {{\rm{e}}^ - } \to {\rm{Ag}}\left( s \right)

On multiplying the cathode reaction by 2, the reaction is written as follows:

2Ag+(aq)+2e2Ag(s){\rm{2A}}{{\rm{g}}^ + }\left( {aq} \right) + 2{{\rm{e}}^ - } \to 2{\rm{Ag}}\left( s \right)

The net cell reaction is written as follows:

Cu(s)+2Ag+(aq)Cu2+(aq)+2Ag(s){\rm{Cu}}\left( s \right) + {\rm{2A}}{{\rm{g}}^ + }\left( {aq} \right) \to {\rm{C}}{{\rm{u}}^{2 + }}\left( {aq} \right) + 2{\rm{Ag}}\left( s \right)

Ans:

The reaction at anode is written as follows:

Cu(s)Cu2+(aq)+2e{\rm{Cu}}\left( s \right) \to {\rm{C}}{{\rm{u}}^{2 + }}\left( {aq} \right) + 2{{\rm{e}}^ - }

The reaction at cathode is written as follows:

Ag+(aq)+eAg(s){\rm{A}}{{\rm{g}}^ + }\left( {aq} \right) + {{\rm{e}}^ - } \to {\rm{Ag}}\left( s \right)

The net cell reaction is written as follows:

Cu(s)+2Ag+(aq)Cu2+(aq)+2Ag(s){\rm{Cu}}\left( s \right) + {\rm{2A}}{{\rm{g}}^ + }\left( {aq} \right) \to {\rm{C}}{{\rm{u}}^{2 + }}\left( {aq} \right) + 2{\rm{Ag}}\left( s \right)